OpenStudy (gavin39):

.What are the explicit equation and domain for an arithmetic sequence with a first term of 6 and a second term of 2? an = 6 − 2(n − 1); all integers where n ≥ 1 an = 6 − 2(n − 1); all integers where n ≥ 0 an = 6 − 4(n − 1); all integers where n ≥ 0 an = 6 − 4(n − 1); all integers where n ≥ 1

OpenStudy (gavin39):

@minoz

OpenStudy (anonymous):

to calculate nth term, we use a+(n-1)d d=secon term- first term u can verify whether n>=0 or 1 by substituting n=0 or 1 and check for first term 6

OpenStudy (gavin39):

Hmm still not 100% with the answer cause these last couple questions on my HW im cluless on

OpenStudy (gavin39):

Needed to be done at midnight that's why im trying to rush and get the answers

OpenStudy (anonymous):

ans is D

OpenStudy (gavin39):

Thank you a lot!!! Could you help me out with 2 more?

OpenStudy (anonymous):

substitute n=1 in all the equations and check for ans 6 and substitute n=2 and check for ans 2 this is the easy way of verification

OpenStudy (anonymous):

sure

OpenStudy (gavin39):

.Given the exponential equation 2x = 128, what is the logarithmic form of the equation in base 10? x = log base 2 of 2, all over log base 2of 128 x = log base 2 of 128, all over log base 2 of 2 x = log base 10 of 2, all over log base 10 of 128 x = log base 10 of 128, all over log base 10 of 2

OpenStudy (anonymous):

is question 2^x or 2x

OpenStudy (gavin39):

2x

OpenStudy (anonymous):

C

OpenStudy (anonymous):

not 100% sure

OpenStudy (gavin39):

not a problem. Here is the last one

OpenStudy (gavin39):

.What are the explicit equation and domain for a geometric sequence with a first term of 5 and a second term of −10? an = 5(−2)n − 1; all integers where n ≥ 1 an = 5(−2)n − 1; all integers where n ≥ 0 an = 5(−15)n − 1; all integers where n ≥ 1 an = 5(−15)n − 1; all integers where n ≥ 0

OpenStudy (anonymous):

for geometric an=a*r^n-1 here r=second term/first term n>=1

OpenStudy (gavin39):

is it A or C because both are n>=1

OpenStudy (anonymous):

r=-2 so ...

OpenStudy (gavin39):

ooo im sorry its A lol

OpenStudy (gavin39):

Thank you again! I really appreciate it

OpenStudy (anonymous):

welcome :)